Find the absolute extrema of f(x, y) = x^2 + y^2 − 2x − 2y + 1 on the set D = {(x, y): 0 ≤ x ≤ 2 , 0 ≤ y ≤ 2 }
Answer #1

fy)= 2² + y ² - 2x - 2y + LD differunale Lille ? ? 10202 lingy Of content f 9X022.1. 0:20 Differentiate again with & treatingDifferentiate Eq? 27 wrky & koeping. a les constant fry= 240.0 fry : b. Sofrom &Q2 (2 plut ft: 0 22-2=0 HEL from Exo , put. 2 - 250 Y=L We have paint (x, y) = (1,1) These valves dies in the set of (2Istoking for et horum Yelise. If at-520 The then flf17) has a mat vore 870 then f(x, y) hay a minm vous So here in this problAll have funchon f(x1y)) x2 + y ² - 2x - 28 + 1 Cet (0,0) f(0,0) = 1 at (0 ) flo, 1) = 0+1-O-2 tl. at (0, 2) f(0, 2) lo.of u-at (211) 4+1-4-2 + f/2, 1) = et (212). f (212) = 4+4-4-4+1 EL por from cufferent. Wobre of fluny) uce see that abr Minm valir